The sum of two numbers is 112. The greater of the two is 32 less than 11 times the lesser number. Find the number?
Answer by checkley71(8403)

x+y=112 or x=112-y
y=11x-32
y=11(112-y)-32
y=1232-11y-32
y+11y=1232-32
12y=1200
y=1200/12
y=100 the greater number.
x=112-100
x=12 the lesser number.
proof:
100=11(112-100)-32
100=11(12)-32
100=132-32
100=100